### v1

# Generating Synthetic Electronic Health Record (EHR) Data: A Review with Benchmarking

Abstract
Objectives: To conduct a scoping review of existing approaches for synthetic Electronic Health Records (EHR) data generation, to benchmark major methods with proposed open-source software, and to offer recommendations for practitioners.
Materials and Methods: We search three academic databases for our scoping review. Methods are benchmarked on open-source EHR datasets, Medical Information Mart for Intensive Care III and IV (MIMIC-III/IV). Seven existing methods covering major categories and two baseline methods are implemented and compared. Evaluation metrics concern data fidelity, downstream utility, privacy protection, and computational cost.
Results: 42 studies are identified and classified into five categories. Seven open-source methods covering all categories are selected, trained on MIMIC-III, and evaluated on MIMIC-III or MIMIC-IV for transportability considerations. Among them, Generative Adversarial Network (GAN)-based methods demonstrate competitive performance in fidelity and utility on MIMIC-III; rule-based methods excel in privacy protection. Similar findings are observed on MIMIC-IV, except that GAN-based methods further outperform the baseline methods in preserving fidelity. A Python package, “SynthEHRella”, is provided to integrate various choices of approaches and evaluation metrics, enabling more streamlined exploration and evaluation of multiple methods, which can be accessed at https://github.com/chenxran/synthEHRella.
Discussion: Method choice is governed by the relative importance of the evaluation metrics in downstream use cases. We provide a decision tree to guide the choice among the benchmarked methods.
Conclusion: GAN-based methods excel when distributional shifts exist between the training and testing populations. Otherwise, CorGAN and MedGAN are most suitable for association modeling and predictive modeling, respectively. Future research should prioritize enhancing fidelity of the synthetic data while controlling privacy exposure, and comprehensive benchmarking of longitudinal or conditional generation methods.
Keywords: Confidentiality, Diffusion Models, Generative Adversarial Network, Generative AI, Phecode, Privacy, Scoping Review, Synthetic EHR, Transformer, Variational Auto-Encoder.
1 Introduction
Electronic Health Records (EHRs) store patients’ health information collected via encounters with health systems in a digital format, consisting of both structured and unstructured data. EHRs are often linked to biorepositories that enable molecular profiling of blood or tissue samples. Such rich repositories that integrate disease phenotypes, multi-omics markers, medical images, medications, and laboratory results have become a critical resource for research in biomedicine [1, 2] and public health [3, 4, 5]. For researchers in medical informatics, bioinformatics, statistics, computer science, and other quantitative fields, EHR-linked biobanks [6, 7, 8] are valuable for empirical evaluations of analytical methods using multi-modal data [9, 10, 2, 11]. However, few real-world EHR data that are publicly available exist due to patient consent requirements and patient privacy considerations. Even when large databases are publicly available for research, when it comes to evaluation of methods through simulation studies, one needs repeated replicates of realistic EHR data.
Synthetic EHR data generation has emerged as a promising solution to unlock enormous research and educational potential of real-world healthcare data while safeguarding data confidentiality. First, high-quality realistic synthetic EHR data enable the development and evaluation of analytical tools with less privacy concerns [12, 13, 14, 15, 16]. Second, for minority subpopulations, where the risk of reidentification is higher, using privacy-preserving synthetic data is ethicaly more defensible  [17, 18, 19]. Third, synthetic EHR generation facilitates experimenting with the size of the real training and testing data over multiple replicates for accuracy and uncertainty assessment of methods [20, 21]. Finally, synthetic data provide an ideal environment for training healthcare professionals, data scientists and other researchers in EHR usage, clinical workflows, and data analysis without the risk of exposing real patient information [22, 23].
With the rapid growth in the literature on synthetic EHR generation (Figure S1), numerous review and discussion papers have focused on either recent methodological developments or on the drawbacks of existing evaluation metrics [24, 25, 26, 27, 28]. However, there is a notable lack of review papers with comprehensive benchmarking of existing methodologies using benchmark datasets. The latest benchmarking paper [29], while making valuable contributions to evaluating Generative Adversarial Network (GAN)-based [30] synthetic EHR generation methods, has two major limitations. First, it evaluated exclusively the methods that are built around GAN, falling short of offering a more comprehensive comparison against alternative approaches in other categories we identified in Section 3.1. Second, the benchmarking was based on closed-source data, limiting the reproducibility and the feasibility for evaluating future methods by other researchers. Finally, with an explosion in latest generative artificial intelligence (GenAI) methods, including large language models (LLMs) and diffusion models, a more comprehensive and updated review, evaluation, and benchmarking is needed.
In this paper, we address this knowledge gap by conducting a scoping review of the extant literature on synthetic EHR generation methods and provide a comprehensive benchmarking of seven current approaches and two baseline approaches. Our study provides a holistic evaluation of synthetic EHR generation methods using multiple evaluation metrics, with open-source datasets to ensure reproducibility and accessibility for future research. We also provide an open-source Python package “SynthEHRella” that can be readily used on-premise for synthetic EHR methods evaluation.
Based on our literature review, we classify existing methods into five categories: Rule-based, GAN-based, Variational Autoencoder (VAE)-based [31], Transformer-based [32], and Diffusion-based methods [33, 34]. We include 1-2 representative methods from each category that have usable publicly-available codes and compare their performances on three critical aspects: fidelity, downstream utility, and privacy [29, 27, 35]. Specifically, fidelity assesses how closely the synthetic data resemble the real EHR data distributions, often in terms of summary features such as the number of visits per patient or the marginal distribution of International Classification of Disease (ICD) codes or PhecodeX [36], an extended version of the Phecode [37], which is built from human-crafted grouping of ICD codes [^1] We will use “phecode” or “phecodeX” interchangeably to refer to the codes in the PhecodeX system.. Downstream utility evaluates the gap between using synthetic data and the real data when performing downstream tasks. Privacy examines the exposure risk of sensitive patient information (e.g., their inclusions in the dataset, their clinical attributes) in synthetic EHR data. In addition to the quality of synthetic data, we also evaluate the computational time cost associated with each method.
Most prior studies focused on using the MIMIC-III dataset [38] for training the data generation engines and for evaluation. We expand our evaluation to include the newly released MIMIC-IV dataset [39]. The inclusion of the MIMIC-IV dataset enables us not only to evaluate the performance of the existing methods on a larger patient cohort with more recent data, but to examine how would the existing methods that have been trained and evaluated on MIMIC-III perform under a different but related population, namely to assess the transportability of the synthetic data generation methods under distributional shifts. Based on our evaluation results, we provide recommendations for practitioners and identify areas where further developments are required.
We summarize the key contributions of this paper in the following:
  1. 1.
    We present a scoping review and comprehensive, multi-pronged benchmarking of synthetic EHR generation methods on open-source datasets, considering fidelity, utility, privacy, and computational cost;
  2. 2.
    We incorporate the latest MIMIC-IV dataset into our benchmarking, in addition to MIMIC-III, to provide an evaluation of existing methods under potential distributional shift in the training and testing populations to assess transportability;
  3. 3.
    Based on our study, we present a decision tree to offer recommendations to practitioners. We identify key gaps in current approaches and indicate potential avenues for future research;
  4. 4.
    We develop and release the SyntEHRella package to facilitate and encourage future benchmarking and reproducibility efforts in synthetic EHR generation.
The rest of the paper is organized as follows. In Section 2, we introduce methods for literature review and inclusion criteria of the methods included in our benchmarking. We then discuss data source in Section 2.2, and introduce the evaluation procedure, including statement of the evaluation problem, settings, and metrics in Section 2.3. Numerical studies on the effect of varying input training dataset sizes or the generated synthetic dataset sizes are presented in Section 2.4. In Section 3, we present our findings. Specifically, we summarize our results for literature review in Section 3.1; we describe the summary statistics of the data used for benchmarking in Section 3.2; we then present our main findings on evaluation and numerical studies in Section 3.3 and Section 3.4, respectively. We introduce the open-source SynthEHRella package in Section 3.5. In Section 4, we interpret our findings, present a decision tree based on our current evaluation results to offer recommendations for practitioners, discuss challenges of existing methods, outline areas for future methods development, and limitations of this work. The paper concludes with remarks in Section 5.
2 Method
2.1 Literature Review
2.1.1 Search Procedure
In the spirit of a scoping review, we conducted a search on May 5th, 2024 to identify published and preprint articles in English that propose methods for generating synthetic EHR data. The search keywords include (a) synthetic EHR, (b) synthetic EHR generation, and (c) synthetic electronic health records. The search was performed across three academic databases: Google Scholar, PubMed, and Semantic Scholar. An additional search verification was conducted on August 30th, 2024 to identify updated literature.
All search results underwent a rigorous screening by the authorship team to exclude irrelevant or unqualified papers. First, studies that did not focus on pipeline development of synthetic EHR generation are removed. Second, papers without an available codebase for reproducing or generating synthetic EHR data were excluded from the consideration of benchmarking. In our final slate, we chose seven distinct methods from each of the five categories that are frequently present in the literature. We also added two baseline methods, one based on directly resampling from the real dataset not intended for privacy-sensitive use but for anchoring the values of the performance metrics, and another based on generating binary phecodes using independent Bernoulli distributions with the mean equal to the marginal phenotype prevalence of each phecode. Detailed descriptions of these methods and the baselines are provided in Section 3.1.
2.2 Benchmarking Datasets
In this work, the evaluation is carried out on two open-source large-scale Intensive Care Unit (ICU)-based and emergency department (ED)-based [^2] ED-based admissions appear only in the MIMIC-IV dataset. benchmark EHR datasets described as follows:
MIMIC-III [38]
MIMIC-III is a large-scale EHR dataset comprising 58,976 unique ICU admissions at Beth Israel Deaconess Medical Center (BIDMC) in Boston, Massachusetts between 2001 and 2012. The dataset consists of mainly adult patients, as well as 7,874 neonates admitted to the Neonatal Intensive Care Unit (NICU). All patient data have been de-identified to ensure patient privacy. The dataset uses International Classification of Disease codes (ICD-9) to record diagnostic events. Many existing studies on synthetic EHR data have focused their evaluations primarily on the presence or count of codes for diagnoses, medications, and procedures within the MIMIC-III dataset.
MIMIC-IV [39]
MIMIC-IV is a recently released, updated version of MIMIC series datasets that features a larger patient cohort, using both version 9 and 10 of ICD as coding system. MIMIC-IV includes 431,231 hospital admissions of patients that were admitted into either the emergency department or ICUs between 2008 and 2019 at BIDMC. Due to the overlap of data collection periods (2008-2012), a subset of patients in MIMIC-IV also appear in MIMIC-III, although it is difficult, if not impossible, to link profiles across datasets because of the de-identification procedures. As a recently-released dataset, few studies have evaluated their synthetic data methods on MIMIC-IV. To the best of our knowledge,  [40, 41, 42] are the only existing works that have assessed their methods on this dataset.
In this study, version 1.4 of MIMIC-III and version 2.2 of MIMIC-IV are used for the comparative evaluation and benchmarking. To highlight the difference from “synthetic data”, we use “real data” to refer to either the MIMIC-III or MIMIC-IV data which we will distinguish when necessary; we use “training data” to refer to the subset of data for training methods which can be real, synthetic data, or their combination. Note that the scenario where the synthetic data can be treated as training data only appear in the predictive utility evaluation described in Section 2.3.
2.3 Evaluation
2.3.1 Statement of Evaluation Task
In this study, we focus on the generation of cross-sectional diagnostic events data, following the capabilities of most publicly-available codebases. ICU-based (and ED-based) EHR datasets, such as MIMIC-III and MIMIC-IV, have a longitudinal nature, where each individual patient may have multiple encounters. To this end, we convert all longitudinal records into a cross-sectional format by aggregating each patient’s records into a single representation. Formally, consider an EHR dataset with cross-sectional format $\mathcal{D}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\}$, where each entry $\mathbf{x}_{i}=(x_{i}^{(1)},\ldots,x_{i}^{(K)})^{\top}$ is a $K$-dimensional binary vector representing the presence/absence of diseases coded by a specific coding system (e.g., ICD-9). Each dimension $x_{i}^{(k)}=1$ if the $k$-th disease is diagnosed at least once in all visits of $i$-th patient, $k=1,\cdots,K$, $i=1,\cdots,N$. The goal of synthetic EHR generation is to produce a synthetic dataset of size $M$, $\mathcal{D}_{syn}=\{\mathbf{x}_{i,syn}\}^{M}_{i=1}$, where each sample in $\mathcal{D}_{syn}$ can be viewed as samples drawn from the same distribution as those in the real dataset of size $N$ denoted as $\mathcal{D}_{real}=\{\mathbf{x}_{i,real}\}^{N}_{i=1}$, while ensuring privacy protection for patients in $\mathcal{D}_{real}$. We represent the datasets by two $N\times K$ and $M\times K$ binary matrices denoted as $\mathbf{X}_{real}$ and $\mathbf{X}_{syn}$, respectively. For this study, we fix $M=50,000$ except that in the numerical studies (see Section 2.4) we vary size of generated samples to investigate their impacts on the quality of synthetic data.
The entire MIMIC-III dataset is used to train generative models if no pre-trained models are available. For a fair evaluation, we adopt PhecodeX 1.0 [36] as the standard coding system for evaluation. Before evaluation, all synthetic EHR data are first mapped to the ICD-10 coding system, and then to PhecodeX. Specifically, we map ICD-9 to ICD-10 based on the icd-mappings package in Python [^3] https://pypi.org/project/icd-mappings/0.1.1/; we map Systematized Medical Nomenclature for Medicine–Clinical Terminology (SNOMED-CT) to ICD-10 via the SNOMED mapping tool [^4] https://prod-mapping.ihtsdotools.org/#/project/records; we then convert ICD-10 coding system to PhecodeX according to the Phecode Map X (Extended) [^5] https://phewascatalog.org/phecode_x. Based on MIMIC-III, we identify $K=2,254$ distinct phecodes. For visualization, we keep only phecodes with more than 50 patients in MIMIC-III, resulting in $K=1,773$ phecodes. For quantitative evaluation, on the other hand, all additional granularity (i.e., digits after the decimal point) in phecodes were omitted and merged into their corresponding parent categories, resulting in $K=595$ parent phecodes.
2.3.2 Performance Evaluation Metrics
We evaluate the quality of synthetic EHR generated by all the selected methods and baselines on both MIMIC-III and MIMIC-IV from three widely adopted perspectives: 1) fidelity for evaluating how closely the distribution of synthetic EHR data matches with real EHR data; 2) downstream utility for measuring the performance gap in downstream analyses compared to real data; 3) privacy exposure for assessing how well the synthetic EHR generation methods protect patient information. Additionally, we also evaluate the selected methods’ computational efficiency. The evaluation metrics are presented below.
Fidelity
To evaluate whether the generative algorithms capture phenotype-level prevalence information, we first evaluate the dimension-wise distributional discrepancy between the real and the synthetic datasets. We compute the Maximum Mean Discrepancy (MMD) to evaluate the absolute difference in prevalence. Similarly, we use Root Mean Squared Percentage Error (RMSPE) and Mean Absolute Percentage Error (MAPE) for assessing the relative difference in prevalence. These metrics are computed as follows:
$\hat{\mu}^{(k)}_{syn}=\frac{1}{M}\sum^{M}_{i=1}x^{(k)}_{i,syn},\leavevmode\nobreak\ \quad\hat{\mu}^{(k)}_{real}=\frac{1}{N}\sum^{N}_{i=1}x^{(k)}_{i,real},\quad k=1,\dots,K;$
$\texttt{MMD}=\max_{1\leq k\leq K}\left|\hat{\mu}^{(k)}_{syn}-\hat{\mu}^{(k)}_{real}\right|;$
$\texttt{RMSPE}=100\times\sqrt{\frac{1}{K}\sum_{i=1}^{K}\left(\frac{\hat{\mu}^{(i)}_{syn}-\hat{\mu}^{(i)}_{real}}{\hat{\mu}^{(i)}_{real}}\right)^{2}};$
$\texttt{MAPE}=\frac{100}{K}\sum_{i=1}^{K}\left|\frac{\hat{\mu}^{(i)}_{syn}-\hat{\mu}^{(i)}_{real}}{\hat{\mu}^{(i)}_{real}}\right|.$
All the RSMPE values are divided by a factor of 100 for ease of presentation.
In addition to quantitative evaluation, we also present visualizations to compare the distribution characteristics between the real and the synthetic EHR data. Specifically, we use a boxplot to show the dimension-wise prevalence using each method, and another boxplot to show the distribution of the number of unique phecodes per patient in the data generated by each method. A scatterplot is included to compare the dimension-wise prevalence between real and synthetic EHR data, with each dot representing a phecode.
We are also interested in evaluating whether generative models preserve pair-wise correlations between the phecodes. To this end, we calculate the Pearson correlation matrix following [29, 43] for both real and synthetic datasets and compute the correlation Frobenius distance (CFD):
$\hat{\Sigma}_{real}={\rm Corr}(\mathbf{X}_{real}),\quad\hat{\Sigma}_{syn}={\rm Corr}(\mathbf{X}_{syn}),$
$\texttt{CFD}=\left\lVert\hat{\Sigma}_{real}-\hat{\Sigma}_{syn}\right\rVert_{F}=\sqrt{\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}(a^{real}_{kk^{\prime}}-a^{syn}_{kk^{\prime}})^{2}},$
where $a^{real}_{kk^{\prime}}$ and $a^{syn}_{kk^{\prime}}$ represents the ($k,k^{\prime}$)-th element in correlation matrices $\hat{\Sigma}_{real}$ and $\hat{\Sigma}_{syn}$, respectively. In addition, we visualize element-wise difference of correlation between the real and the synthetic data via boxplots; 100,000 pairs of entries from the upper triangle of the two correlation matrices are sampled and shown in the boxplots.
In addition, considering that each disease dimension consists of binary data indicating occurrence, we present another metric based on the Frobenius distance between co-occurrence matrices for the real and the synthetic datasets. The co-occurrence matrices Frobenius distance (COFD) is calculated as follows:
$\mathbf{B}_{real}=\mathbf{X}^{\top}_{real}\mathbf{X}_{real},\quad\mathbf{B}_{syn}=\mathbf{X}^{\top}_{syn}\mathbf{X}_{syn},$
$\texttt{COFD}=\left\lVert\mathbf{B}_{real}-\mathbf{B}_{syn}\right\rVert_{F}=\sqrt{\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}(b^{real}_{kk^{\prime}}-b^{syn}_{kk^{\prime}})^{2}},$
where $b^{real}_{kk^{\prime}}$ indicates the number of patients having both the $k$-th and $k^{\prime}$-th phenotypes in the real data; similarly, $b^{syn}_{kk^{\prime}}$ for the synthetic data. All the COFD values are divided by a factor of 1000 for ease of presentation.
Finally, we assess whether we can distinguish the synthetic and the real data used to train the synthetic data generation method, referred to as discriminative prediction. We use both the real and the synthetic data to train a logistic regression model for predicting whether a given sample in the combined data is real or synthetic. We validate the model via 5-fold cross-validation. We report the Area Under the Curve (denoted as AUC) and Accuracy (denoted as ACC) of the classifier, where lower values indicate higher resemblance of the synthetic data to the real data.
Utility
A primary objective of generating synthetic EHR data is to replace the use of real data in downstream applications, thereby protecting patient privacy without compromising data utility. To this end, we evaluate the utility of synthetic data by examining their performance in both analytical and predictive tasks.
Firstly, analytical utility assesses the capability of the synthetic data in providing inferential results that approximate those produced based on the real data. To this end, we focus on the analytical task of estimating associations between binary phenotypes $x_{i}^{(k)}$ and $x_{i}^{(k^{\prime})}$ by fitting a logistic regression model with the $k$-th phenotype as the outcome, and the $k^{\prime}$-th phenotype as the explanatory variable. The point estimates and the 95% confidence intervals of regression coefficients (i.e., log-odds ratios) of the predictor phenotype are reported. A larger overlap in the 95% Confidence Intervals (CIs) between those obtained based on the real data and the synthetic data indicates higher analytical utility of the synthetic data generation method.
Secondly, predictive utility measures the capability of the synthetic data in training machine learning (ML) methods for downstream use. To evaluate existing methods’ predictive utility, a binary classification task is constructed, where the $k$-th phenotype $x^{(k)}_{i}$ is treated as the outcome with the remaining phenotypes in $\mathbf{x}_{i}$ as predictors. In this work, we list three scenarios commonly used in the literature [29, 44] that concern different combinations of the sources of data used for training and testing a ML method. All the testing are based on the real data for objective evaluations. Suppose we split the real dataset into training and testing subsets (4:1 ratio). All the three scenarios are visualized in Figure S2, with detailed descriptions as follows:
  1. 1.
    “Train on Synthetic, Test on Real” (TSTR). To mimic common scenarios in practice where synthetic data are used, we train an ML model using the synthetic data and test the method on the real data. This setting evaluates whether the synthetic data can substitute the real data in downstream predictive tasks;
  2. 2.
    “Train on Synthetic + Real, Test on Real” (TSRTR). This setting investigates if the synthetic data can be considered as a data augmentation of the real data in downstream applications which have been shown to be capable of boosting performance by preventing overfitting issues when only a limited amount of real data is available [45] We stack the training split in the real data and the synthetic data together to train the ML model and test it on the test set of the real data. In practice, this is only feasible when the investigator has access to the real data. Nonetheless, it is included here for showing whether synthetic data has deleterious or beneficial effect when added to a real data;
  3. 3.
    “Train on Real and Test on Real”(TRTR). This scenario does not use synthetic data. The ML model is trained on the training data split, and find the test error on the testing data split. This provides the baseline in predictive utility evaluation.
For each of the above three scenarios, we calculate the AUC and ACC of the classifiers. Because TRTR is a real data based approach, we use it as a baseline and report the differences: ${\rm AUC}_{TSTR}-{\rm AUC}_{TRTR}$ and ${\rm AUC}_{TSRTR}-{\rm AUC}_{TRTR}$, and similarly for ACC. Higher values indicate better utility of the synthetic data.
In this study, analytical utility evaluations are conducted under two tasks. In the first task, whether a patient gets any type of cancer (phecodeX starting with “CA”) is treated as the outcome, and Obesity (EM_236) is treated as the predictor; for the second task, Diabete (EM_202) is treated as the outcome, and Hypertension (CV_401) is treated as the predictor. See Table 2 for prevalences of any type of cancers, obesity, hypertension, and diabetes. For predictive utility evaluation, Hypertension (CV_401) is treated as the outcome, because it is the most prevalent phenotype in both MIMIC-III and MIMIC-IV cohorts; the remaining phenotypes are treated as predictors. The logistic regression model is treated as our ML models, with results of more advanced ML methods, including Random Forest [46], Gradient Boosting [47], XGBoost [48], LightGBM [49], and K-Nearest Neighbours (K-NN), reported in Supplementary Materials.
Privacy
First, we use membership inference risk (MIR) [50] to assess the risk that an attacker, who gains access to a patient’s complete medical record, could determine whether this patient is included in the training dataset. A subject’s membership to the training dataset may be sensitive because the patient may wish to keep this information private from potential attackers. For example, the training data may be based on all the HIV-positive patients treated in a regional facility. Latest work of membership inference attack using synthetic data proposed to conduct privacy attack via an attacker model that can detect local overfitting which occurs if the synthetic data density evaluated at the real data record has a high value; this typically occurs when the real data record is close to the bulk of synthetic data [51]. We compute the minimum Euclidean distance between each real medical record and the synthetic EHR dataset:
$d_{i}=\texttt{min}_{\mathbf{x}\in\mathcal{D}_{syn}}\texttt{dist}(\mathbf{x}_{i,real},\mathbf{x}),$
where $\mathbf{x}_{i,real}$ is a record from the real dataset. We further provide an average characterization by reporting the mean and median of these minimum distances as the dataset-level risk of membership inference if all the subjects in the real data may be attacked:
$\texttt{MIR}_{mean}=\frac{1}{N}\sum^{N}_{i=1}d_{i},\leavevmode\nobreak\ \texttt{MIR}_{median}=\texttt{median}(d_{1},\dots,d_{N}),$
where higher values indicate lower risks of membership inference attack. Note that unlike [29, 43] where $d_{i}$ is also calculated, we do not further specify a threshold to classify into successful or failed attacks and calculate the corresponding F1 score to evaluate the membership inference risk of the synthetic data, because the evaluation results may be sensitive to the subjective threshold value.
The scenario differs from the discriminitive prediction in the fidelity evaluation in that 1) unlike the use of both the real and the synthetic data, membership inference uses only the synthetic data and a real medical record of the subject being attacked; 2) unlike the binary label of real or synthetic data, membership inference concerns the binary label of whether the subject being attacked is a member in the real data curated by the facility; and 3) unlike the AUC and ACC metric for discriminative prediction, membership inference risk here is evaluated by how close is the real medical record of the attacked individual to the synthetic data.
When $\mathbf{x}_{i,real}$ belongs to MIMIC-III, $d_{i}$ measures the risk of identifying the membership of subject $i$ in MIMIC-III based on the synthetic data. When $\mathbf{x}_{i,real}$ belongs to MIMIC-IV, $d_{i}$ measures the risk of identifying whether subject $i$ is a member of MIMIC-III which remains meaningful because MIMIC-III and MIMIC-IV datasets overlapped in the periods of data collection (2008-2012), hence containing overlapping subjects. The $\texttt{MIR}_{mean}$ and $\texttt{MIR}_{median}$ then provide an average characterization of this distance assuming all the patients in MIMIC-IV may subject to membership inference attacks.
In addition, we assess attribute inference risk (AIR) [52] which occurs when attackers who have access to partial information about a patient’s real medical records could infer the missing attributes based on the synthetic data. To evaluate AIR, we select the top 10 balanced (prevalence closest to 0.5) and top 10 imbalanced diseases as unknown attributes simultaneously to be inferred. We applied a 1-Nearest Neighbor (1-NN) approach to match each real record with the most similar record in the synthetic dataset based on the remaining known attributes. The corresponding features in the matched synthetic records are considered as the prediction of the missing attributes in the real records. The F1 score is reported to measure how well the masked attributes can be predicted, where a lower value indicates a lower risk of attribute inference attack.
2.3.3 Computational Cost
Beyond the above three metrics for assessing the quality of the synthetic data, the computational cost of generating synthetic EHR datasets is also an important factor during methods selection. This is especially true if a large amount of synthetic data are needed for evaluating frequentist performance of analytic and prediction methods. To this end, we evaluate the time needed for all the selected methods in synthesizing 100 EHR samples. The evaluations are conducted on a multi-node cluster equipped with Intel(R) Xeon(R) Gold 6148 CPUs with a total of 40 physical cores. A 16GB Tesla V100 GPU is used for running deep generative models when necessary.
2.4 Numerical Studies: Effectiveness of Varying $N$ and $M$
Importantly, we explore the effect by varying the size of training ($N$) and synthetic ($M$) data on the quality of synthetic EHR data. We address two questions. First, given a fixed amount of training data, how many synthetic samples can a generative model produce before the benefits of adding one additional synthetic sample diminish? Although deep generative models such as GANs, VAEs, and diffusion models can theoretically generate infinite samples, the amount of novel information does not increase proportionally with the size of the synthetic dataset, as it is constrained by the finite training data. Second, how does the size of the training data impact the quality of the generated synthetic data, given a fixed number of synthetic samples?
To explore these questions, we select the overall best-performing methods from our benchmarking study. For the first question, we train the selected method and generate $M=1K,2K,5K,10K,20K,50K,100K,200K,500K$ synthetic samples. We evaluate the generated data using three metrics: MMD for fidelity, TSTR for utility (AUC is reported), and AIR for privacy. For the second question, we train the selected method with $N=1K,2K,5K,10K,15K,20K,25K,30K,35K,40K$ real samples and generate $50K$ synthetic samples; we evaluate them with the same metrics. The results are visualized to illustrate how these metrics evolve under different values of $M$ and $N$.
3 Results
3.1 Summary of Literature Review
Our literature search identified 107 publications based on keyword searches, from which 42 studies are found to propose methodologies for generating synthetic EHRs that are relevant to our work. Table 1 lists all these relevant studies, organized by the year of publication. Based on the literature review, we classify most existing methods for synthetic EHR generation into five primary categories:
  • Rule-based method: These approaches generate synthetic EHR data using predefined algorithms and rules, often incorporating real-world disease prevalence and national census data as hyperparameters to simulate disease occurrence and population demographics. Rule-based methods were the dominant approach to generate synthetic EHRs prior to the emergence of deep generative models. From our review, 6 out of 42 methods are rule-based methods [53, 54, 55, 56, 57].
  • GAN-based method: With the rise of deep learning, GANs have become widely adopted for synthetic EHR generation. [58] and [59] are among the first to introduce GANs for this purpose, with subsequent studies focusing on improving the capability of GAN-based models and addressing privacy concerns. Based on the scoping review, 18 out of 42 methods are GAN-based methods [60, 61, 22, 62, 17, 63, 64, 65, 66, 67, 68, 44, 69, 42, 70].
  • VAE-based method: As another strand of deep generative modeling that are popular in the artificial intelligence (AI) community, VAEs are applied to learn latent embedding from the real EHR datasets and generate synthetic EHR data based on this learned structure [18]. Additionally, [41] extends VAEs for longitudinal EHR data generation. Our review shows that 3 out of 42 methods apply VAEs to generate synthetic EHR data.
  • Transformer-based method: Transformers [32], known for their ability to process sequential data [71] and their success in natural language processing [72, 73], have also been explored for longitudinal synthetic EHR generation. In our literature review, 4 out of 42 methods learn to generate synthetic EHR data with the transformer structures. [74] are the first to train transformers on real EHR data for this purpose. Moreover, [75] explores the potential of Large Language Models (LLMs) trained on natural texts in synthesizing EHR data.
  • Diffusion-based method: Most recently, diffusion models [33, 34], which have demonstrated strong performance in image and video generation [76, 77, 78], have been applied for synthetic EHR generation. 8 out of the 42 studies have proposed diffusion-based models for this purpose [79, 80, 81, 82, 43, 83, 84, 85].
In addition to the five main categories above, conventional statistical methods were also used to generate synthetic EHR data. For example, [86] proposed an algorithm based on Gibbs sampling to generate synthetic EHR data; [87] introduced Bayesian Network (BN) in synthetic EHR generation. Furthermore, other neural network structures were also explored in synthetic EHR generation [88].
Based on these literature, we identify three key functionalities that are critical for the design of synthetic data generation algorithms:
Generation conditional on specific variables:
Supporting synthetic conditional EHR generation, which produces data based on specific contexts, is critical for downstream applications that require subgroup analyses or the inclusion of socio-demographic factors. Our literature review reveals that 24 of the 42 studies cover conditional generation. For example,  [59] generates synthetic data based on specific labels to augment downstream predictive applications;  [17] generates data conditioning on specific socio-demographic concepts to overcome small data size challenges in training;  [18] generates synthetic data given specific medical conditions of interest for further scientific investigation.
Generation of repeated visits over time:
As EHR datasets include patients with multiple visits in nature, it is important to capture the longitudinal patterns to generate realistic synthetic timped-stamped data. Of the 42 reviewed studies, 23 incorporate temporal information for longitudinal generation. For example, Synthea [89] generates synthetic EHR longitudinally by generating synthetic patients and simulate their encounters across their lifespans; TimeDiff [68] and PromptEHR [74], on the other hand, applied neural networks with architectures designed to model sequential data to simulate longitudinal EHR.
Generation of multiple modalities of data:
EHR datasets are inherently multimodal, containing various types of medical encounters such as diagnoses, medications, procedures, chart events, clinical notes, and laboratory results. Understanding which modalities each method can generate is crucial for effective data generation. In this study, we focus on events that could be characterized by coding systems (i.e., diagnoses, medications, procedures) and summarize if existing methods could model those events. As shown in Table 1, 16 of the 18 existing studies that evaluate their methods on MIMIC-III primarily focus on diagnostic events; [60, 64, 41, 83, 75] additionally include medication or procedures in their modeling; [58, 74, 90, 85] consider all the three types of events in their modeling.
3.1.1 Methods Selected for Benchmarking
Despite the recent development of synthetic EHR methods, 25 out of 42 studies remain closed-source at the time of this scoping review, as highlighted in Table 1. The lack of accessible codebases poses challenges for fair comparisons across methods and limits the broader use of synthetic EHRs in downstream applications. To conduct benchmarking, we select seven generation methods that provide publicly available and user-friendly codebases. These methods cover the five primary categories discussed above. Specifically, for rule-based methods, we select Plasmode [54] and Synthea [89]; for GAN-based methods, we choose MedGAN [58] and CorGAN [64]; for VAE-based methods, we use a baseline implemented by CorGAN’s authors due to the lack of easily reproducible codebases in existing VAE-based methods; for transformer-based methods, we select PromptEHR [74]; and for diffusion-based methods, we choose EHRDiff [43]. For all the existing methods, we follow their default settings, except necessary minor changes for debugging.
In addition to the methods selected from existing literature, we include two simple baselines for comparison. The first baseline, which we refer to as Prevalence-based Random, randomly generates synthetic EHR data based on the estimated marginal phenotype prevalences without taking phenotype correlation into account. For the second baseline, we construct synthetic EHR data by simply bootstrapping the real EHR dataset; we refer to this method as Resample. Note that while the Resample baseline is impractical for synthetic data generation when privacy concern exists, it serves as a valuable reference by providing an upper bound of the fidelity and utility capability of the existing synthetic EHR generation methods, as it closely resemble the real data. To account for possible loss of information during code transformation (Section 2.3.1), we compute phenotype prevalence based on the MIMIC-III data in ICD-9 system for Prevalence-based Random to generate synthetic data. Similarly, we bootstrap real data under the same coding system for Resample.
3.2 Summary Statistics of Benchmark Datasets
Table 2 presents the summary statistics for the MIMIC-III and MIMIC-IV datasets. Overall, the socio-demographic distributions are largely similar between MIMIC-III and MIMIC-IV, except that MIMIC-III additionally includes 7,874 neonates in the dataset. More than half of the patients in MIMIC-III have been diagnosed with cancer (52.9%), higher than the prevalence in MIMIC-IV (40.2%). The prevalence of mental health conditions, such as anxiety and depression, is higher in MIMIC-IV (13.0% and 20.8%, respectively) compared to MIMIC-III (11.0% and 16.0%, respectively). Whereas diabetes and hypertension are more common in the MIMIC-III cohort (25.4% and 47.6%, respectively) compared to the MIMIC-IV cohort (21.7% and 46.0%, respectively). The prevalence of obesity, on the other hand, is higher in the MIMIC-IV cohort (12.6%) than in the MIMIC-III cohort (5.4%). Regarding EHR characteristics, both MIMIC-III and MIMIC-IV exhibit right-skewed distributions for the number of encounters per person, unique phecodes per person, and length of follow-up. The median numbers of encounters per person for both datasets are 1.0, indicating that over half of patients in both data visited ICU or ED (in MIMIC-IV) only once during the data collection period. The mean number of encounters is higher in MIMIC-IV (2.4 visits) than that in MIMIC-III (1.3 visits). In terms of the length of follow-up, significant zero-inflation issues are observed for both datasets, with a median of 0 year of follow-up. This is consistent with our previous findings that half of patients have only one hospital admission during data collection period. The average length of follow-up, on the other hand, is higher in MIMIC-IV (1.2 years) than that in MIMIC-III (0.3 years). The average number of unique phecodes per patient is greater in MIMIC-III (31.3 phecodes) than in MIMIC-IV (26.0 phecodes), although MIMIC-IV has a higher maximum number of unique phecodes for a single patient (260 phecodes) than MIMIC-III (203 phecodes).
3.3 Evaluation Results
In this subsection, we present the evaluation results of the selected methods and baselines. By default, we use the entire MIMIC-III dataset to train synthetic data generation methods. We evaluate the quality of the synthetic data by comparing against the real MIMIC-III data or against MIMIC-IV for transportability considerations.
3.3.1 Fidelity
Figure 1 shows the boxplots of the marginal prevalences in the real and the synthetic EHR data. The Prevalence-based Random baseline, along with MedGAN, CorGAN, and EHRDiff effectively preserve the distribution information of marginal prevalence of MIMIC-III. In contrast, Plasmode, Synthea, and VAE tend to consistently underestimate the prevalence compared to the real dataset. In addition, the MIMIC-IV dataset exhibits a similar distribution in the estimated prevalence compared to MIMIC-III, with a median of the prevalences across all phenotypes slightly lower than that in MIMIC-III.
Figure 2 displays the number of unique phecodes per patient in EHR data. CorGAN shows similar distributions to MIMIC-III, whereas most of the remaining methods tend to underestimate the number of unique phecodes per patient. Conversely, PromptEHR consistently overestimate the number of unique phecodes, and Prevalence-based Random shows higher values at 25th and 50th quantiles in the number of phecodes per patient compared to MIMIC-III.
Figure S3 compares the phecode-wise prevalence between the synthetic data and the real data (MIMIC-III). Prevalence-based Random achieves an almost perfect match with the real data but tends to overestimate the prevalence for several non-rare phecodes, which may due to code transformation. Among the selected methods, MedGAN and CorGAN exhibit less divergence from the real data, while Plasmode, PromptEHR, VAE, and EHRDiff show greater discrepancies. Notably, Synthea generates non-zero prevalences for only 160 phecodes, likely due to its limited support for only around 90 disease modules, and potential information losses during multi-step code transformations from the SNOMED-CT [91] to the PhecodeX coding system.
On the capability of preserving bivariate relationships, Figure 3 shows the boxplots of the element-wise differences in the correlation matrix. Most of the methods, except MedGAN which underestimates the correlations, overestimate the associations. This issue is even more pronounced when compared against MIMIC-IV, where all the methods consistently overestimate correlations. MedGAN, however, continues to underestimate correlations.
Table 3 presents the quantitative evaluation of dimension-wise and correlation-wise discrepancy, as well as discriminative prediction results. Prevalence-based Random consistently outperform the selected methods across most fidelity metrics, except for the Prevalence-based Random baseline’s higher CFD score of 46.9, which is expected since it does not account for bivariate correlations by design. Among the selected methods, CorGAN performs best in preserving distributional information, especially with its strong performance in CFD attributable to a module designed for modeling associations. Similar results are observed for the COFD that assesses co-occurrence rate discrepancies.
When evaluated on the MIMIC-IV data, CorGAN outperforms the two baseline approaches in preserving the phecode-wise prevalence, achieving an MMD of 0.234 for absolute differences. However, PromptEHR demonstrates the best performance in the relative difference metrics (RMSPE: 113.5; MAPE: 1418.7), likely caused by a better capacity of modeling rare diseases. Additionally, because the synthetic data were generated by training on MIMIC-III, when compared against MIMIC-IV, the difficulty of distinguishing the synthetic from the real MIMIC-IV data is reduced significantly. Among them, CorGAN emerge as the relatively more effective method, with AUC score of 0.927. In terms of the COFD, PromptEHR and MedGAN, instead, achieve the best results, with COFD 887.76 and 1017.83, respectively.
3.3.2 Utility
We first focus on the analytical utility of the synthetic EHR data. Table 4 presents the logistic regression results. In the Obesity vs. All Cancers log odds ratio estimation, the estimated coefficients for real data are 0.87 (95% CI: [0.78, 0.96]) in MIMIC-III and 0.81 (95% CI: [0.78, 0.84]) in MIMIC-IV. In the Diabetes vs. Hypertension log odds ratio estimation, the estimated log odds ratio in the MIMIC-III and MIMIC-IV datasets are 1.22 (95% CI: [1.17, 1.26]) and 1.41 (95% CI: [1.39, 1.44]), respectively. Although their directions are consistent across datasets, the magnitudes differ from each other in the second task. For the synthetic data, Prevalence-based Random yields insignificant estimates for both tasks. MedGAN, CorGAN, PromptEHR, and EHRDiff all produce estimated log-odds ratio with direction consistent with that obtained from the MIMIC-III real data. Notably, CorGAN has estimated log-odds ratio of 1.08 (95% CI: [0.97, 1.19]) and 1.17 (95% CI: [1.12, 1.21]) in the two tasks, with confidence intervals substantially overlapping those of MIMIC-III in the second task. On the other hand, PromptEHR and EHRDiff consistently underestimate the log-odds ratio compared to the real data, with 0.51 (95% CI: [0.42, 0.61]) and 0.57 (95% CI: [0.53, 0.60]) in the first task, and 0.60 (95% CI: [0.55, 0.64]) and 0.95 (95% CI: [0.89, 1.00]) in the second task, respectively. Notably, VAE encounters positivity issues in the first task, where it fails to generate patients with obesity and absent any type of cancer, leading to failure in convergence during model fitting. This issue highlights the limitations of VAE in generating reliable synthetic data.
We then examine the predictive utility of synthetic EHR data, with results presented in Table 3. For TSTR, all the selected methods show performance degradation when compared to TRTR on both MIMIC-III and MIMIC-IV. Among selected methods, MedGAN shows the smallest performance drop in AUC on both datasets, with a decrease of 0.063 on MIMIC-III and 0.076 on MIMIC-IV. In contrast, Plasmode and EHRDiff exhibit the most significant declines, with AUC decreases of 0.268 and 0.178, respectively, on MIMIC-III, and 0.218 for both on MIMIC-IV.
In the TSRTR evaluation, on MIMIC-III, Plasmode becomes competitive with MedGAN, showing a minor AUC improvement of 0.0003 and an ACC improvement of 0.003, respectively. When evaluated using the MIMIC-IV data, MedGAN, VAE, PromptEHR perform competitively, all bringing a marginal improvement of 0.004 in ACC for the downstream task.
3.3.3 Privacy
The capacities of the selected methods in privacy protection are presented in Table 3. Prevalence-based Random and Synthea demonstrate the strongest performance in preventing membership inference attacks on MIMIC-III, largely because these methods do not rely on training data to generate synthetic EHRs.
In terms of preventing attribute inference attacks, the rule-based methods, Synthea and Plasmode, perform best, with predictive performance by attackers at 0.571 and 0.595, respectively. Prevalence-based Random follows closely, with a predictive performance of 0.622. Plasmode’s superior performance can be attributed to its dimension-by-dimension data generation approach, which omits the associations in the real dataset.
When evaluated on the MIMIC-IV dataset (assuming subjects in MIMIC-IV can be attacked), rule-based methods (i.e., Plasmode and Synthea) achieve the best results in lowering the risk of both membership and attribute inference attacks, mostly consistent with findings if MIMIC-III subjects are attacked.
3.3.4 Computational Cost
We now turn to the computational costs of the selected methods, with results presented in Table S1. Overall, rule-based methods have higher computational costs compared to most of the deep generative methods in other categories. Specifically, Synthea and Plasmode cost 24 and 12,292 seconds to generate 100 samples, respectively [^6] Although Plasmode is extremely expensive in computation under this setting, the cost increases only sub-linearly with sample sizes. For example, generating 50,000 synthetic samples with Plasmode costs 17,225 seconds..On the other hand, among deep generative methods, GAN-based and VAE-based methods are significantly faster than Transformer-based and Diffusion-based approaches. Specifically, CorGAN and MedGAN generate 100 samples in 0.26 and 0.08 seconds, respectively, while VAE requires 0.16 seconds. In contrast, EHRDiff takes 0.65 seconds to generate the same number of samples, and PromptEHR achieves the highest computational cost among deep generative methods at 77 seconds per 100 samples, which is primarily because of the longitudinal nature of its EHR generation process.
3.4 Numerical Studies
Based on the evaluation results shown in Figure 123 and Table 3 and 4, CorGAN demonstrates outstanding performance across all three key perspectives of synthetic data quality. Hence, CorGAN is selected for the numerical studies to explore the effect of varying sizes of $M$ and $N$ on the quality of the synthetic data.
Effect of varying synthetic data size $M$:
Figure 4 presents the results of varying the synthetic sample sizes ($M$) on the performance of synthetic EHR data. Fidelity improves (the metric MMD declines) consistently as the number of generated samples increases, which is expected since sufficiently large sample sizes are required to approximate the real data distribution. When the number of synthetic samples exceeds 10,000, the MMD score begins to fluctuate and stabilizes between 0.048-0.049. In terms of utility, increasing the synthetic sample size narrows the performance gap between synthetic and real data, although the rate of improvement is non-linear. For privacy, larger synthetic sample sizes increases the risk of the attribute inference attack, indicating that generating more synthetic EHR data increases the risk of information leakage from the training dataset.
Effect of varying training data size $N$:
Figure 5 illustrates the results of varying training data sizes on synthetic EHR data performance. As the training data size increases, both the fidelity and utility of the generated synthetic data improve. However, this improvement comes with an increased risk of privacy exposure, indicating a trade-off between synthetic data quality and privacy protection.
3.5 SynthEHRella: A Package for Benchmarking Synthetic EHR Performance
One of the major contributions of this paper is to consolidate all of the 7 synthetic data generation pipelines under one unified coding framework. To this end, we develop a Python package, “SynthEHRella”, to integrate various choices of approaches and evaluation metrics, enabling more streamlined exploration and evaluation of multiple methods before deciding on a particular choice in applications. This reduces the barrier of comparing multiple methods, making it an extremely useful benchmarking tool for developing new methods. The package is available at https://github.com/chenxran/synthEHRella.
4 Discussion
In this paper, our scoping review and benchmarking provide several key insights. First, Prevalence-based Random is the best-performing method for resembling the real MIMIC-III data, though it, by design, does not capture bivariate associations. This finding, which indicates a major area of improvement in data fidelity, is consistent with previous work [29]. Second, CorGAN has a superior performance among the selected methods in preserving data distributions and analytical utility, particularly in capturing bivariate associations. It even outperforms the baselines in resembling the MIMIC-IV cohort. However, CorGAN consistently underestimates the prevalences compared to the real data as shown in Figure S4. Third, as expected, Synthea offers significant advantages in privacy protection, because it does not rely on training data. However, its utility is limited by the restricted number of disease generation modules available. Fourth, GAN-based methods are the most computationally efficient in generating synthetic data compared to methods in other categories, likely due to their lightweight design and the exclusion of longitudinal EHR data in most implementations. Lastly, summary statistics and fidelity visualization reveal a significant distributional discrepancy between MIMIC-III and MIMIC-IV, making them well-suited for evaluating the transportability of the synthetic EHR methods.
The evaluation results reveal a trade-off between fidelity, utility, and privacy protection in synthetic EHR data generation. To further explore these relationships, we present a heatmap of the correlations among the evaluation metrics for both datasets in Figure S7. For consistency, all “lower-is-better” metrics are inverted so that positive correlations indicate synergistic relationships and negative correlations indicate trade-offs. In MIMIC-III, we observe negative associations in all the utility-privacy metric pairs, as well as the fidelity-privacy pairs, suggesting that improved utility and fidelity may come at the expense of privacy exposure. When evaluated on MIMIC-IV, negative associations are observed between the privacy and almost all the non-privacy evaluation metrics, which are consistent with the findings on MIMIC-III. These findings suggest that the trade-offs observed above are common despite distributional differences between the MIMIC-III and the MIMIC-IV data.
In addition, we also explore the relationships between the evaluation metrics across the two datasets. As shown in Figure S8, for each metric, there are strong positive correlations between MIMIC-III and MIMIC-IV, suggesting that transportability of the existing methods is closely related to their performance in the distribution they are trained on. In addition, negative associations are observed between privacy and all the remaining non-privacy metrics, showing that privacy-related trade-offs remain across datasets.
Based on our evaluation, we offer recommendations for practitioners to select synthetic EHR generation methods as shown in Figure 6. In scenarios where there are distributional shifts between training and testing samples, CorGAN is the preferred method except when downstream predictive performance is the primary objective, in which case MedGAN is the best choice. When the training and testing populations are identical, we recommend MedGAN for downstream predictive tasks, CorGAN and Prevalence-based Random for better estimation of dimension-wise prevalence, and CorGAN for better estimation of associations among the phenotypes.
Our evaluation identifies several open opportunities. First, there is significant room for enhanced or new methods that improve on fidelity. We showed that marginal Prevalence-based Random can generate synthetic data that resemble the real data, despite failing to capture correlations. However, even CorGAN, the best in fidelity in our evaluation, consistently overestimated these prevalences, leaving gaps for improvements. Second, for specific sub-populations defined by socio-demographic variables, it is unclear whether training the synthetic EHR generation methods on the entire population is the most effective strategy and needs further studies. For example, MIMIC-III contains a significant number of neonatal patients with distinct distributional characteristics than the adult patients. Scientific studies often either excluded adult [92, 93] or neonatal and pediatric patients [94, 95, 96] to conduct analysis. However, synthetic data are often generated by methods trained on all patients combined. The consequence of this mismatch between the population used to generate synthetic data and the analytic subpopulation is less well understood and needs further research. Lastly, as the privacy concerns may vary significantly depending on the context of use, it will be useful to develop a framework and tools to fix privacy exposure risk at a pre-specified level and then optimize fidelity or utility.
This work has several limitations. First, the benchmarking focuses exclusively on diagnostic events, excluding other important modalities such as procedures, medications, clinical notes, and medical imaging. We also did not focus on the evaluation of longitudinal synthetic EHR data. As modeling and evaluation under these settings would require additional considerable effort, we leave it as future work. Furthermore, as we hope this work could create a unified platform for evaluating and inspire further development and benchmarking of synthetic EHR methods, we developed SynthEHRella to allow for the extension of evaluations to other modalities. Second, with the rapid expansion of this area, this study does not cover all the available methods for benchmarking. However, our publicly available evaluation package facilitates the addition of new methods to the benchmark. Third, when evaluating synthetic EHR generation on MIMIC-IV, we excluded diagnostic events not present in MIMIC-III to ensure consistent evaluation dimensions, which may limit insights into domain adaptation. Future research should develop better evaluation metrics for domain adaptation scenarios. Fourth, while we selected key evaluation metrics, the study does not cover all the metrics ever used in previous works. However, we believe that our selected metrics address the most critical concerns. Our benchmarking package also enables the inclusion of additional metrics as needed. Fifth, the benchmarking in this study did not create a composite weighted evaluation metrics to compare the methods across different scenarios. Assigning weights to different metrics can be application-dependent. We leave the development of such a metric as a potential future direction. Sixth, as the decision tree is obtained based on our evaluation results of the seven selected methods and baselines, it is not necessarily always applicable when we face different EHR datasets or populations with different distributions from the MIMIC-III/IV cohorts. We suggest similar evaluation pipelines in these settings, e.g., by adapting the SynthEHRella package which is released for public use. Seventh, the analytical utility evaluation in this work, conducted via bivariate logistic regressions, overlaps conceptually with the evaluation of correlation difference in fidelity evaluation. This overlap is primarily because the lack of socio-demographic characteristics in the generated EHR data. However, analytical utility evaluation provides further comparisons of the confidence intervals on the association of popular diseases. Furthermore, in the future one could add confounders from methods capable of generating data with socio-demographic information to evaluate their ability in preserving adjusted associations. Eighth, this study focuses only on the transportability of models trained on MIMIC-III to MIMIC-IV, without considering the possibility of training models directly on MIMIC-IV. While extending existing methods to train on a subset of MIMIC-IV data is possible, we leave this for future work. Lastly, in this work we assumed that no missing data in the real EHR data  [97, 44]. Future benchmarking work is needed to evaluate the implication of missing data in the quality of synthetic EHR generation.
5 Conclusion
This paper presents a scoping review and a benchmarking of the existing representative methods with open-source codebases for generating synthetic EHR data using benchmark datasets. Under privacy concern, when distributional discrepancy exists between the training and testing samples, CorGAN is the best choice for dimension-wise prevalence estimation and downstream analytical tasks; MedGAN is recommended for training ML models for downstream use. Otherwise, when the distribution of the testing population is similar to that of the training data, MedGAN, Prevalence-based Random, and CorGAN are preferred choices for predictive modeling, dimension-wise prevalence estimation, and modeling associations among phenotypes, respectively. Despite the significant increase in the number of publications in the recent literature on synthetic EHR generation methods, we identify a few critical areas for future research. The benchmarking results highlight a substantial performance gap in fidelity between the existing methods and Prevalence-based Random, indicating the need of future methods research to narrow this gap. Finally, comprehensive benchmarking of the existing synthetic data generation methods that account for the longitudinal nature of EHR data and support conditional data generation is needed.
Acknowledgements
This work was supported through grant DMS1712933 from the National Science Foundation and MI-CARES grant 1UG3CA267907 from the National Cancer Institute. The funders had no role in the design of the study; collection, analysis, or interpretation of the data; writing of the report; or the decision to submit the manuscript for publication. Conflicts of Interest: None.
Table 1: A summary of existing methods for synthesizing EHR data, outlining the functionalities included, aspects of evaluation included, data source used, the availability of code, and the event modality covered in each method. Abbreviation: DWS: dimension-wise similarity; CMS: correlation matrix similarity; DP: discriminative prediction; Human: Human Evaluation; TSTR: Train on Synthetic, Test on Real ; TSRTR: Train on Synthetic + Real, Test on Real; MIR: Membership Inference Risk; AIR: Attribute Inference Risk. †: this method only evaluated on MIMIC-IV. ✓  denotes that the respective method includes the listed functionality, evaluation metric, data source, or event modality; ✗; indicates the lack of inclusion of that aspect. The “-” symbol is used in the event modality columns when the MIMIC-III dataset is not used as a data source by the method.
References Model Type Year Conditional Longitudinal Fidelity Utility Privacy Open-Source Use MIMIC-III Event Modality
DWS CMS DP Human TSTR TSRTR MIR AIR Diagnosis Procedure Medication
[53] EMERGE Rule 2010 - - -
[86] PeGS non-param 2013 - - -
[54] Plasmode Rule 2014 - - -
[55] PADARSER/GRiSER Rule 2014 - - -
[56] CoMSER Rule 2016 - - -
[59] RGA/RCGAN GAN 2017 - - -
[58] MedGAN GAN 2017
[89] Synthea Rule 2018 - - -
[57] ATEN Rule 2018 - - -
[60] medBGAN/medWGAN GAN 2019
[61] GcGAN GAN 2019 - - -
[22] HealthGAN GAN 2020
[62] DAAE GAN 2020
[17] EMR-WGAN/EMR-CWGAN GAN 2020 - - -
[63] HGAN GAN 2020 - - -
[64] CorGAN GAN 2020
[65] ADS-GAN GAN 2020 - - -
[66] SMOOTH-GAN GAN 2020 - - -
[87] Bayesian Network Bayesian Network 2021
[67] SynTEG GAN 2021 - - -
[18] EVA VAE 2021 - - -
[74] PromptEHR Transformer 2022
[68] LongGAN GAN 2022 - - -
[40] Graph-VAE VAE (Graph) 2023 - - -
[44] EHR-M-GAN GAN 2023 - - -
[79] EHR-TabDDPM Diffusion 2023
[80] MedDiff Diffusion 2023
[88] LSTM-EHR LSTM 2023 - - -
[81] ScoEHR Diffusion 2023
[90] HALO Transformer 2023
[82] TimeDiff Diffusion 2023
[97] EHR-Safe GAN 2023
[43] EHRDiff Diffusion 2024
[41] MSIC VAE 2024
[83] EHR-D3PM Diffusion 2024
[98] CEHR-GPT Transformer 2024 - - -
[84] FlexGen-EHR Diffusion 2024 - - -
[69] Bt-GAN GAN 2024 - - -
[42] IGAMT GAN 2024   ✓ - - -
[75] LLM-EHR Transformer/LLM 2024
[70] LDP-GAN GAN 2024 - - -
[85] EHRPD Diffusion 2024
Table 2: Summary statistics of the MIMIC-III and MIMIC-IV data. For age, gender, ethnicity, and diseases, the values represent percentages; counts are shown in parentheses.
MIMIC-III MIMIC-IV
# of patients 46,520 180,733
# of ICU patients 46,520 50,920
# of Admissions 58,976 431,231
Demographics
[…]
   0-9 16.9 (7,874) 0.0 (0)
   10-19 0.8 (371) 1.8 (3,210)
   20-29 3.6 (1,672) 12.6 (22,711)
   30-39 4.4 (2,029) 11.6 (20,937)
   40-49 8.7 (4,040) 12.2 (21,965)
   50-59 14.2 (6,586) 16.7 (30,167)
   60-69 16.9 (7,845) 17.4 (31,474)
   70-79 17.0 (7,918) 14.0 (25,323)
   80+ 17.6 (8,185) 13.8 (24,946)
[…]
   Female 43.9 (20,399) 53.0 (95,729)
   Male 56.2 (26,121) 47.0 (85,004)
[…]
   Asian 3.6 (1,690) 4.2 (7,583)
   Black/African American 8.3 (3,864) 13.0 (23,538)
   Hispanic 3.5 (1,642) 5.4 (9,766)
   White 69.6 (32,372) 67.2 (121,421)
   Other/Unknown 14.9 (6,952) 10.2 (18,425)
[…]
Anxiety 11.0 (5,112) 13.0 (23,412)
Cancer 52.9 (24,620) 40.2 (72,683)
Depression 16.0 (7,431) 20.8 (37,551)
Diabetes 25.4 (11,811) 21.7 (39.211)
Hypertension 47.6 (22,147) 46.0 (83,107)
Obesity 5.4 (2,518) 12.6 (22,855)
EHR Characteristics
Encounter per person
   Minimum 1.0 1.0
   Median 1.0 1.0
   Mean 1.3 2.4
   Maximum 42.0 238.0
Unique phecode per person
   Minimum 0.0 0.0
   Median 26.0 14.0
   Mean 31.3 26.0
   Maximum 203.0 260.0
Length of follow-up (years)
   Minimum 0.0 0.0
   Median 0.0 0.0
   Mean 0.3 1.2
   Maximum 11.5 12.0
Figure 1: Boxplots of the estimated phecode-wise prevalences (Y-axis) across distinct phecodes in the real datasets or synthetic datasets generated from the selected methods shown on the X-axis. Only the $K=1,773$ distinct phecodes occurring in more than 50 patients in MIMIC-III are included.
Figure 2: Boxplots of the number of unique phecodes per patient (Y-axis) based on the real datasets or synthetic datasets generated from selected methods shown on X-axis. Only the $K=1,773$ distinct phecodes occurring in more than 50 patients in MIMIC-III are included.
(a) MIMIC-III.
(b) MIMIC-IV.
Figure 3: Boxplots of the differences in the estimated Pearson correlation for $100,000$ randomly selected pairs of phecodes (out of all (${1,773\choose 2}$ pairs) comparing the ones obtained from synthetic data with those from (a) MIMIC-III and (b) MIMIC-IV for each of the 9 selected methods. The solid red line represents an exact match in correlation; the area within the red dotted lines indicates a correlation difference of less than 0.01.
Table 3: Quantitative evaluation results on synthetic data quality when evaluated using the MIMIC-III data (top block) or MIMIC-IV data (bottom block). For fidelity, Maximum Mean Discrepancy (MMD), Root Mean Squared Percentage Error (RMSPE), Mean Absolute Percentage Error (MAPE), Correlation Frobenius Distance (CFD), Co-occurrence matrix Frobenius distance (COFD), and the Area Under the Curve (AUC) and Accuracy (ACC) of discriminative prediction are reported. For utility, the performance gap of AUC and ACC between Train on Synthetic, Test on Real (TSTR) and Train on Real, Test on Real (TRTR) are reported, and similarly for the Train on Synthetic + Real, Test on Real (TSRTR) one. For privacy, the mean and median of minimum Euclidean distance between each real medical record and the synthetic EHR dataset are reported for Membership Inference Risk (MIR); the F1 score of predictive performance of 1-Nearest Neighbor (1-NN) are reported for Attribute Inference Risk (AIR). $\downarrow$  represents the lower the better; $\uparrow$  represents the higher the better. For each metric, the overall top two methods are in black boldface; the best method excluding Resample and Prevalence-based Random is in blue boldface. Red boldface indicates the method is in overall top-two and also ranks the best among all but Resample and Prevalence-based Random.
Fidelity Utility Privacy
MMD RMSPE MAPE CFD COFD AUC ACC TSTR (AUC) TSTR (ACC) TSRTR (AUC) TSRTR (ACC) MIR (mean) MIR (median) AIR
$\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$
MIMIC-III
Resample 0.0040 0.17 7.18 4.42 29.50 0.476 0.500 0.007 0.005 0.008 0.006 0.97 0.00 0.952
Prevalence-based Random 0.0044 0.14 6.99 46.94 114.11 0.494 0.421 -0.287 -0.275 -0.014 -0.010 5.47 5.29 0.622
Synthea 0.7206 11.56 222.96 47.34 469.45 0.999 0.995 -0.149 -0.115 -0.004 -0.002 4.86 4.80 0.571
Plasmode 0.5945 127.41 1951.77 39.01 596.81 0.999 0.996 -0.268 -0.333 0.0003 0.001 4.53 4.58 0.595
MedGAN 0.1430 0.58 46.85 120.89 252.11 0.913 0.840 -0.063 -0.038 -0.003 0.003 2.93 2.83 0.874
CorGAN 0.0478 0.57 44.30 22.20 84.48 0.720 0.669 -0.078 -0.056 -0.020 -0.013 3.06 3.16 0.838
VAE 0.3083 0.94 92.67 44.05 345.86 0.984 0.943 -0.086 -0.063 -0.010 -0.002 3.88 4.00 0.845
PromptEHR 0.8593 2.47 91.31 24.03 762.25 0.990 0.957 -0.081 -0.049 -0.014 -0.011 3.48 3.46 0.839
EHRDiff 0.3381 18.30 245.55 31.98 337.31 0.995 0.967 -0.178 -0.216 -0.029 -0.024 3.66 3.61 0.737
MIMIC-IV
Resample 0.2363 1047.70 11849.43 30.15 1103.51 0.898 0.854 -0.066 -0.028 -0.0009 0.005 2.58 2.00 0.766
Prevalence-based Random 0.3740 1092.77 12540.96 44.92 1153.67 0.956 0.914 -0.254 -0.168 -0.005 -0.001 2.96 2.45 0.694
Synthea 0.7216 727.39 5689.49 58.61 1491.42 0.999 0.993 -0.127 -0.092 -0.001 -0.002 3.80 3.46 0.597
Plasmode 0.5970 5844.53 84377.27 53.23 1535.78 0.999 0.999 -0.218 -0.206 $0$0.0002 3.98 3.46 0.497
MedGAN 0.2377 797.89 9259.80 107.59 1017.83 0.962 0.916 -0.076 -0.059 -0.0007 0.004 2.83 2.45 0.714
CorGAN 0.2335 959.54 10549.75 38.92 1152.84 0.927 0.883 -0.131 -0.081 -0.004 0.001 2.86 2.45 0.732
VAE 0.2360 680.68 6566.97 53.89 1430.53 0.978 0.923 -0.123 -0.074 -0.004 0.004 3.44 3.16 0.765
PromptEHR 0.8686 113.46 1418.71 39.09 887.76 0.995 0.980 -0.117 -0.066 -0.003 0.004 3.09 2.83 0.750
EHRDiff 0.3176 1274.05 14207.65 44.11 1309.36 0.995 0.981 -0.218 -0.152 -0.007 0.0008 3.04 2.65 0.675
Table 4: Estimated $\beta$ from the real data and synthetic data generated by selected methods in two bivariate logistic regression models: 1) Obesity (EM_236) v.s. with any type of cnacers (all phecodeX starting with “CA”); 2) Diabetes (EM_202) v.s. Hypertension (CV_401). Synthetic data generated by VAE fails to converge for the first task as no patients with obesity and absent any type of cancer are generated.
Obesity v.s. All Cancers Diabetes v.s. Hypertension
$\hat{\beta}$ 95% CI $\hat{\beta}$ 95% CI
Real Data MIMIC-III 0.871 (0.782 0.960) 1.219 (1.174 1.263)
MIMIC-IV 0.810 (0.782 0.838) 1.412 (1.388 1.437)
Synthetic Data Resample 0.868 (0.781 0.955) 1.194 (1.151 1.237)
Prevalence-based Random -0.125 (-0.316 0.066) 0.003 (-0.038 0.043)
Synthea 1.685 (1.614 1.756) 3.590 (3.512 3.668)
Plasmode -0.039 (-0.143 0.066) 0.027 (-0.118 0.171)
MedGAN 3.636 (3.453 3.819) 0.864 (0.822 0.905)
CorGAN 1.080 (0.970 1.191) 1.165 (1.120 1.209)
VAE - - 3.190 (3.063 3.318)
PromptEHR 0.514 (0.420 0.608) 0.597 (0.553 0.642)
EHRDiff 0.565 (0.527 0.603) 0.946 (0.891 1.000)
Figure 4: The performance under varying sizes of the synthetic data ($M$) generated by CorGAN trained on full real data. The chosen evaluation metrics for the three perspectives are Maximum Mean Discrepancy (MMD), Train on Synthetic, Test on Real (TSTR) measured by Area Under the Curve (AUC), and F1 score for Attribute Inference Risk (AIR), respectively. The synthetic data generation process is repeated five times, and the mean evaluation metrics are reported. Note that for MMD and F1 score in AIR, lower values indicate better performance.
Figure 5: The performance of the synthetic data generated by CorGAN trained under varying sizes of the real data for training ($N$). The chosen evaluation metrics for the three perspectives are Maximum Mean Discrepancy (MMD), Train on Synthetic, Test on Real (TSTR) measured by Area Under the Curve (AUC), and F1 score for Attribute Inference Risk (AIR), respectively. The synthetic data generation process is repeated five times, and the mean evaluation metrics are reported. Note that for MMD and F1 score in AIR, lower values indicate better performance.
Figure 6: Decision tree for selecting synthetic EHR generation methods to generate phenotype EHR data as informed by our evaluation results of the representative methods studied in this paper. The numbers in parentheses indicate the computational time cost of corresponding methods for generating 100 samples.
Appendix S1 Supplementary Materials
Figure S1: The number of publications on synthetic EHR generation methods in recent years.
Figure S2: Three different training approaches described in predictive utility. Real data (either MIMIC-III or MIMIC-IV) are split into training and testing subsets with 4:1 ratio. For TRTR, ML models are trained on the training subset of the real data, and are tested on the real testing subset. For TSTR, ML models are trained on all the synthetic data, and tested on the real testing subset. For TSRTR, the real training subset is stacked with all the synthetic data to train ML models, and the testing subset of the real data is used to evaluate the trained ML models.
Figure S3: Scatterplots comparing the phecode-wise prevalences estimated from the MIMIC-IV data or synthetic data with the one from the real data in the MIMIC-III dataset. The red dashed lines indicate the perfect agreement of estimated prevalence between the synthetic and the MIMIC-III real data.
(a) MIMIC-III.
(b) MIMIC-IV.
Figure S4: Boxplots of the ratio of the estimated PhecodeX-wise prevalence between the real and the synthetic data generated by CorGAN, grouped by PhecodeX categories defined in [36]. Y-axis represents the ratio of estimated prevalence, less than 1 means underestimation by the synthetic data. X-axis represents all the PhecodeX categories. Red line indicates the median of prevalence ratio of all phecodes. The significant over-estimation within the neonatal phenotype group in the MIMIC-IV dataset is because of its lack of neonatal population.
(a) Evaluated on MIMIC-III.
(b) Evaluated on MIMIC-IV.
Figure S5: The difference of AUC between TSTR and TRTR based on different ML methods trained on either MIMIC-III or MIMIC-IV. The large performance gap in Resample is because of the fact that the TRTR baseline is trained on the training subset of the MIMIC-III/IV dataset. On the other hand, Resample bootstrap over the entire MIMIC-III/IV datasets to “synthesize” data.
(a) Evaluated on MIMIC-III.
(b) Evaluated on MIMIC-IV.
Figure S6: The difference of AUC between TSRTR and TRTR based on different ML methods trained on either MIMIC-III or MIMIC-IV. The large performance gap in Resample is because of the fact that the TRTR baseline is trained on the training subset of the MIMIC-III/IV dataset. On the other hand, Resample bootstrap over the entire MIMIC-III/IV datasets to “synthesize” data.
Table S1: Computational time cost of different synthetic EHR generation methods. Time is recorded for generating 100 samples. †: the generation time of Plasmode does not increase linearly with size of generation (e.g., Plasmode spends 17,225 seconds in generating 50,000 synthetic samples). The top-2 fastest results are bolded.
Computational cost (sec/100 samples)
PromptEHR 71.0
MedGAN 0.08
CorGAN 0.26
VAE 0.16
Synthea 24.0
Plasmode $12291.5^{\dagger}$
EHRDiff 0.65
(a) Evaluated on MIMIC-III.
(b) Evaluated on MIMIC-IV.
Figure S7: Within-dataset Pearson correlation among different evaluation metrics in Table 3 for (a) MIMIC-III, and (b) MIMIC-IV, respectively. All “lower-is-better” evaluation metrics (e.g., Maximum Mean Discrepancy (MMD)) are flipped to their opposite numbers before calculating the Pearson correlation. Positive correlations (grids in red) indicate synergistic relationships between two metrics, and negative correlations (girds in blue) indicate trade-offs. No numbers are shown on the diagonal because here we focus on the relationships between distinct metrics evaluated on the same data; only lower triangular numbers are shown because of symmetry.
Figure S8: Cross-dataset Pearson correlations of the evaluation metrics (in Table 3) between MIMIC-III and MIMIC-IV. All “lower-is-better” evaluation metrics (e.g., Maximum Mean Discrepancy (MMD)) are flipped to their opposite numbers before calculating the Pearson correlation. Positive correlations (grids in red) indicate synergistic relationships between two metrics, and negative correlations (girds in blue) indicate trade-offs. We need the numbers on the diagonal and the upper triangular matrix because here we focus on the relationships of each metric evaluated on two distinct data sets, MIMIC-III and IV.
S1.1 Details of Generating Synthetic EHR Data from Existing Methods
This subsection introduces the procedures for generating synthetic EHR data using existing methods. Specifically, for MedGAN, CorGAN, and VAE, and EHRDiff, we train the models from scratch using the publicly released code by their authors with the entire MIMIC-III dataset, and then generate the synthetic data. For PromptEHR, we use the pretrained checkpoint released by the authors to generate synthetic data. For Synthea, we generate synthetic data based on the default setting. In the case of Plasmode, which generates only one outcome variable at a time, the synthetic EHR data are generated dimension by dimension. For each dimension, the presence of this disease is treated as the outcome variable, and the time variable is defined as either the time to the first diagnosis of the disease or the time to the patient’s last ICU visit, depending on the presence of disease. The treatment variable is randomly sampled from a Bernoulli distribution with probability 0.5. Key socio-demographic characteristics, including age, gender, and ethnicity, are treated as covariates to fit the Cox Proportional Hazards model.